MY 11. [-17.14 Points] DETAILS LARCALCET7 8.7.040. Use a table of integrals to find the indefinite integral. (Use for the con use a table of integrals to find the indefinite integral for botb
Answer #1

✓ [cat (o)) do Integral Reduction Cotnt (a) Lcot Cada : - ſcot -2(a) da n- Hue n = 4 Cot col s Cat odo: Coat ce [ Cat cos2 -da 19-24 Let u=x2 duzzada xdx= 1 de du del . 259-12 Fue du Let Hi 3 sincul du= 3 Coscu) du (9-u² = 19-9 sin rcu) Jack-bint